4. Find the inverse Laplace transform of 2s + 24 F(S) = (s +1)(s? +65 +34) Solution:
Answer #1

2 st 24 F(S) - 25+24 (st)(8+66+34) (st)[(s +37+8 ] 2st 24 А sti + Bsto (stil (1s +35757] (st35t 5 = 25+24= A(5+65 +34) + (sty